OpenStudy (anonymous):

Right triangle h=.12m l= 3.5 m find sin(theta)

OpenStudy (anonymous):

.12^2 + 3.5^2 = 12.2644 Sqrt 12.2644 = 3.502 sin theta = .12/3.502 = .0342 .0342 * (180/pi) = 1.96 degreees Correct?

OpenStudy (anonymous):

Yup

Directrix (directrix):

@Operative The posted answer does not look correct to me. From the given information, how do we know where angle theta is located. Same question for h and l.